OpenStudy (anonymous):

simplify each rational expression. x^2+10x+25/ x^2+9x+20

OpenStudy (anonymous):

I always had problemsfactoring, Can you help me.

OpenStudy (anonymous):

how do you get that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2+10x+25/ x^2+9x+20 (x+5)(x+5) / (x+5)(x-4) = x+5/ x - 4 restriction: x cannot be = 4

OpenStudy (anonymous):

@Ket-kat can 5 be x?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

well, lets plug in five and see what happens

OpenStudy (anonymous):

Oh, So if it =0 then x can't be that number.

OpenStudy (anonymous):

x+5/x-4 f(5) = (5) + 5/ (5) - 4 = 10/1 =10

OpenStudy (anonymous):

The denominator (the bottom) can NEVER be equal to zero. So whenever you have a fraction, simplify the whole equation; and then set the bottom to equal to zero in order to find your zeros like so: x+ 5/x-4 x-4=0 x=4 whatever this number is, x CANNOT BE

OpenStudy (anonymous):

but the top can be 0 though, right

OpenStudy (anonymous):

Yes, unless the problem you're working says otherwise

OpenStudy (anonymous):

If the top is = 0, then the answer will ALWAYS be : UNDEFINED

OpenStudy (lgbasallote):

it's x+4 not minus

OpenStudy (anonymous):

so i did it all worng.

Parth (parthkohli):

If the bottom is = 0 not top

Parth (parthkohli):

@Ket-kat you messed it up

OpenStudy (anonymous):

oops, I'm only human xD young and messy lol

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (anonymous):

You didn't do it wrong

OpenStudy (anonymous):

so the answer would be x doesnt = -4 and -5

OpenStudy (anonymous):

well, if you're factoring and then solving for zeros you should be okay (it would be + 4)